### v1

# Causal Predictive Optimization and Generation for Business AI

Abstract.
The sales process involves sales functions converting leads or opportunities to customers and selling more products to existing customers. The optimization of the sales process thus is key to success of any B2B business. In this work, we introduce a principled approach to sales optimization and business AI, namely the Causal Predictive Optimization and Generation, which includes three layers: 1) prediction layer with causal ML 2) optimization layer with constraint optimization and contextual bandit 3) serving layer with Generative AI and feedback-loop for system enhancement. We detail the implementation and deployment of the system in LinkedIn, showcasing significant wins over legacy systems and sharing learning and insight broadly applicable to this field.
Causal Machine Learning, Mixed Integer Programming, Explainability, Generative Artificial Intelligence , Sales Process Optimization
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: XXXX; XXX, 2025; XXXXisbn: 978-1-4503-XXXX-X/18/06ccs: Mathematics of computing Probability and statisticsccs: Computing methodologies Natural language processingccs: Computing methodologies Machine learningccs: Applied computing Enterprise computingccs: Applied computing Decision analysis
1. Introduction
Sales functions play a critical role in success and revenue growth of any Business-to-Business (B2B) and Software-as-a-Service (SaaS) business. Each member of the sales fleet manages anywhere from a few to hundreds of accounts, and their main objective is to convert leads and upsell to existing customers. Sales Reps spend majority of their time figuring out the right account or person to talk to and what to talk about. Determining the optimality of each component with these accounts on hand is a highly time-consuming task. An efficient business AI engine will greatly improve productivity of sales fleet and is key to the optimization of the sale process and success of the business.
More specifically, addressing the following challenges that exist in the sales ecosystem is critical.
  • Inefficient allocation of sales resources - Absence of causal modeling and insights lead to wasted efforts on low impact accounts while missing high-value opportunities.
  • Complex trade-offs in decision-making – Balancing multiple objectives and constraints in sales (e.g., revenue growth, engagement, and resource limits) is difficult without a structured optimization approach.
  • Lack of trust and adaptability in recommendations – Sales teams struggle with black-box like model outputs, and static systems fail to incorporate real-time feedback, reducing long-term effectiveness.
To address these challenges, we propose Casual Predictive Optimization and Generation (CPOG) as an end-to-end framework solution.
1.1. Related Work
The application of Machine Learning (ML) in sales has advanced considerably, as companies increasingly harness data to optimize decision-making, enhance customer engagement, and streamline operations.
ML models play a crucial role in account prioritization by guiding sales teams toward high-potential prospects based on behavioral patterns and engagement signals. These models also improve sales forecasting by leveraging both structured data  (Rohaan et al., 2022) and textual data  (Elalem et al., 2023), and uplift modeling  (Gutierrez and Gérardy, 2017) enables more accurate predictions. Additionally, customer segmentation through clustering techniques facilitates better sales planning by identifying key audience groups for targeted strategies  (Gautam and Kumar, 2022). B2B customer churn prediction (De Caigny et al., 2021), customer conversion in e-commerce (Gubela et al., 2019).
In recent years, causal machine learning, which combines the traditional causal inference and modern machine learning, enables predictions of individual heterogeneous effect of future potential action (Rubin, 1978)(Kennedy, 2023)(Van der Laan, 2006)(Gutierrez and Gérardy, 2017). Multiple researches shows the great potential of causal ML application to the industry(Shi et al., 2024; Tang et al., 2023). Optimization frameworks transform predictive scores into actionable recommendations and some recent examples in the literature include neural optimization with adaptive heuristics (NOAH) for an intelligent marketing system (Wei et al., 2024) and data-driven sales optimization with regression and chaotic pattern search (Gaddam et al., 2024). In addition, contextual bandit systems have been proven to provide personalized recommendations (Ban et al., 2024).
The rise of explainable AI (XAI) and generative AI (GAI) is transforming how businesses adopt ML models by making complex sales predictions more interpretable  (Bohanec et al., 2017) and interaction of sales engine more seamless. Model-agnostic, rule-based explanations help interpret the results  (Rajapaksha and Bergmeir, 2022), model explanation methods have been shown to provide narratives and explanations for consumers of model outputs.
There is very limited work yet on end-to-end solutions that enable sales functions to harness the power of ML, and provide combination of Prediction, Optimization and Explanation at large scale. Previously mentioned optimization frameworks developed by (Wei et al., 2024) and (Gaddam et al., 2024) cannot be applied to sales account prioritization optimization because they differ in domain ((Wei et al., 2024) was designed for intelligent marketing) and objective ((Gaddam et al., 2024) focused on lead generation). As far as we know, CPOG is the first sales optimization framework focused on account prioritization. The contributions the CPOG framework are discussed in the next section.
1.2. Our Contribution
To our knowledge, CPOG is the first end-to-end framework solution for sales optimization in business AI. In addition, CPOG is the first deployed sales optimization engine in industry setting that
  • utilize causal machine learning to measure the incremental impact of sales actions in the prediction layer, which avoid the inefficient use of sales resources;
  • has dedicated optimization layer to make trade-off of multiple different objective and constraints in sales process, and recommend sales action that balances exploration and exploitation;
  • serves recommendation with explainability, GAI components and feedback-loop, forming a cohesive human-in-the-loop systems that works seamlessly with sales teams.
2. System Overview
CPOG addresses the challenges of the sales ecosystem through a structured AI-driven framework that optimizes sales decisions. The architecture comprises multiple layers: a Data Foundation layer that utilizes engagement, spending, and rep-interaction data; a Prediction Layer that applies causal machine learning models to estimate long- and short-term impact given sales rep intervention; an Optimization Layer that incorporates business constraints and prioritization logic using mixed-integer programming and bandit algorithms; and a Serving Layer that delivers recommendations with human-interpretable explanations and feedback loops.
In addition, the serving layer integrates Generative AI (GAI) to improve user experience and explainability, providing detailed intuitive insights that increase trust in AI-driven recommendations. By incorporating a feedback loop for continuous refinement, the CPOG framework performs as a cohesive human-in-the-loop system, seamlessly aligning with sales teams to optimize decision-making and maximize revenue impact. Overall, this AI-driven system streamlines account prioritization, personalizes recommendations, and improves sales productivity, ultimately driving business growth. In the subsequent sections, we provide a detailed discussion of each layer and its components.
Figure 1. Overall CPOG Architecture
[…]
Overall CPOG Architecture
3. Prediction Layer
The prediction layer predicts a group of performance metrics which sales functions can influence through sales actions. Denoting the feature matrix for customers as $x$, feature matrix of possible sales actions as $x^{a}$ and the performance metrics as $y$, the prediction layer builds a functional mapping:
$y=f(x,x^{a})$
We consider two types of metrics: (1) monetization metrics with longer feedback periods like revenue, customer loyalty and customer LTV, (2) customer engagement metrics like product utilization and product adoption with shorter feedback periods. The goal of the prediction layer is estimate the impact of sales actions on monetization metrics and predict near term customer engagement metrics. The functional mapping $f(.)$ can be learned from historical data on customer outcomes.
3.1. Monetization Uplift Models
Monetization uplift models refer to the set of casual ML models predicting the incremental impact of sale action on monetization metrics. Since sales functions have limited human resources, it is crucial that these resources are put on highest incremental impact. Conventional approach of directly predicting the monetization metrics and recommending accounts with highest ranks can result in approaching accounts that will convert without sales outreach and thus waste of sales resources.
We follow the potential outcome framework (Rubin, 1974) in causal inference where $Y$ denotes the desired outcome of certain monetization metric, $a$ denotes the treatment s.t. $a\in\{1,0\}$ and $x^{U}$ are the pre-treatment variables e.g product utilization, revenue before the treatment etc.
(1) $Y_{a}=f(x^{U}|a)$
We predict $Y_{a}$ using ML models trained on different subsets of the data i.e. $f(.)$ is trained on data from customers in the control or treatment group. The nature of the ML models used depends on the quantity of data - a simple regression model will suffice for smaller datasets, tree-based models for medium sized data and neural networks for large and complex datasets. The uplift from sales outreach is estimated as follows:
$y^{U}=Y_{1}-Y_{0}.$
Meta-learners (Athey and Imbens, 2016)(Künzel et al., 2019) are discrete treatment CATE estimators that model typically on two or more surfaces $Y_{t}$. Meta-learners provide the flexibility of selecting both the estimation model used at each stage along with cross-validation for more data-adaptive estimations at each stage. (Künzel et al., 2019) coined the terms ”S-learner”, ”T-learner” and ”X-learner” which are some common meta-algorithms for estimating heterogeneous treatment effect.
3.2. Customer Engagement Forecast Models
The second objective of the prediction layer is estimating customer engagement metrics, $E$, like product utilization and product adoption.
$\begin{split}y^{E}=f(x^{E})\end{split}$
Where $x^{E}$ are customer engagement features like prior customer engagement metrics, customer firmographics etc. If a product has low data volume e.g. a newly launched product, simple regression models will suffice. As more data and signals become available, more advanced regression models like gradient boosted trees or neural networks might be better suited.
4. Optimization Layer
The optimization layer within the system adeptly transforms predictive scores into actionable recommendations for sales representatives, employing a dual-strategy approach that synergizes both constrained (convex) optimization and contextual bandit methodologies. The constrained optimization component is designed to harmonize key business objectives, including maximizing monetization uplift and enhancing engagement metrics, all while adhering to predefined operational constraints. This component methodically determines the optimal pairings of accounts and sales representatives, ensuring alignment with strategic business goals.
In parallel, the contextual bandit component augments this framework by dynamically adjusting the daily actions assigned to each sales representative. This adjustment mechanism is acutely focused on immediate, actionable items tailored to specific business needs, such as boosting engagement, preventing customer churn, and promoting upsells. Together, these components work in concert to refine the system’s response to evolving market conditions and representative interactions.
4.1. Constrained Optimization
4.1.1. Formulating the Constrained Optimization
Let $\mathcal{C}=\{c_{1},c_{2},\dots,c_{N}\}$ represent the set of customer accounts and =$\mathcal{R}=\{r_{1},r_{2},\dots,r_{M}\}$ the set of sales representatives. Define the binary decision variable $a_{i,j}\in\{0,1\}$ to indicate whether account $c_{i}$ is assigned to representative $r_{j}$. The objective is to maximize a weighted combination of monetization uplift ($U_{i}$) and engagement improvement ($E_{i}$), where the weights dynamically adjust.
(2) $\max_{a_{i,j}}\sum_{i=1}^{N}\sum_{j=1}^{M}\Big{[}w(d_{i})\cdot y^{U}_{i}+(1-w(d_{i}))\cdot y^{E}_{i}\Big{]}\cdot a_{i,j},$
where:
  • $y^{U}_{i}$: Monetization uplift for account $a_{i}$ (normalized to range 0 -100).
  • $y^{E}_{i}$: A function that represents Engagement difference for account $a_{i}$. e.g.
    (3) $y^{E}_{i}=\max\left(|\Delta y^{E1}_{i}|,|\Delta y^{E2}_{i}|,\ldots,|\Delta y^{Ek}_{i}|\right),$
    where $\Delta y^{Ej}_{i}$ denotes the change in the $j$-th engagement metric for the $i$-th account, and $k$ is the total number of engagement metrics considered. (normalized to range 0 -100).
  • $d_{i}$: Days until RTCD (Renwal Target Close Date) for account $a_{i}$. [^1] RTCD (Renewal Target Close Date) refers to the date by which a renewal decision for an account is targeted to be concluded. It is a critical date for sales reps, guiding prioritization and resource allocation to ensure timely engagement and renewal success.
  • $w(d_{i})$: A non-linear Weighting function for $d_{i}$. In this example, it is defined as:
    (4) $w(d_{i})=\frac{1}{1+e^{-k(d_{i}-d_{0})}},$
    where $k$ controls the steepness of the curve and $d_{0}$ centers it.
4.1.2. Constraints
The optimization adheres to the following constraints:
  1. (1)
    Capacity Constraint: Each sales representative $r_{j}$ manages between $n_{\text{min}}$ and $n_{\text{max}}$ accounts:
    (5) $n_{\text{min}}\leq\sum_{i=1}^{N}a_{i,j}\leq n_{\text{max}},\quad\forall i\in\mathcal{C},\forall j\in\mathcal{R}.$
  2. (2)
    Assignment Constraint: Each account is assigned to exactly one representative:
    (6) $\sum_{j=1}^{M}a_{i,j}=1,\quad\forall i\in\mathcal{C},\forall j\in\mathcal{R}.$
  3. (3)
    Eligibility Filters: Accounts must satisfy predefined thresholds:
    • $y^{U}_{i}>T_{U}$: Monetization uplift exceeds threshold $T_{U}$.
    • $y^{E}_{i}>T_{E}$: Engagement improvement exceeds threshold $T_{E}$.
    • For any account $c_{i}$ that has been assigned ($a_{ij}=1$) within the last 14 days, ensure it is not assigned again:
      $a_{ij}=0,\quad\forall j\in\mathcal{R},\text{ if }\sum_{t=T-14}^{T-1}z_{i,t}>0,\quad\forall i\in\mathcal{C}$
      Where
      $z_{i,t}=\begin{cases}1&\text{if }a_{ij}=1,\exists j\in\mathcal{R}\text{ at time t}\\ 0&\text{otherwise}\end{cases}$
  4. (4)
    RTCD Priority: Monetization uplift is prioritized for accounts with $d_{i}$ close to zero, while engagement is emphasized for large $d_{i}$, which is contolled by as $w(d_{i})$ in (2).
4.1.3. Recommendation Rules
This problem is formulated as a Mixed Integer Programming (MIP) problem, which is computationally expensive. To improve tractability:
  1. (1)
    Relaxation: The binary variable $a_{i,j}$ is relaxed to continuous values in $[0,1]$, transforming the problem into a Linear Programming (LP) problem.
  2. (2)
    Solver: The relaxed LP is solved using SCIP (Perron and Furnon, 2024; Bolusani et al., 2024).
  3. (3)
    Post-Processing: Continuous solutions are rounded to binary values, ensuring capacity constraints are respected.
4.1.4. Implementation Notes
  • Weight Calibration: The weighting function $w(d_{i})=\frac{1}{1+e^{-k(d_{i}-d_{0})}}$ balances monetization and engagement based on days to RTCD. We fix $d_{0}=X\times 30$ days (domain knowledge) and tune sharpness $k$ via backtesting to align recommendations with renewal outcomes. This offers interpretability through a fixed anchor and empirical flexibility.
  • Account Matching: $a_{i,j}$ is the decision variable indicating the assignment of accounts $c_{i}$ to sales representatives $r_{j}$.
  • Action Recommendation: Heuristic algorithms are utilized to decide initial action recommendation on the matched account, based on values of $y^{U}_{i}$ and $y^{E}_{i}$. (see Algorithm  1) This serves as a cold start for contextual bandit ranking in section  4.2.
  • Solver Strategy: We relax the MIP and solve the LP via SCIP (Perron and Furnon, 2024; Bolusani et al., 2024) for efficient offline inference. Though MIP is supported, relaxation ensures speed and stability. We plan to explore ML-augmented MIP solvers (recent work by Bertsimas et al. (Bertsimas et al., 2022)) for future low-latency use cases.
Algorithm 1 Decision Making Process for Action Recommendation
  for each account $a_{i}$ do
     Calculate $MonetizationValue_{i}=w(d_{i})\cdot y^{U}_{i}$
     Calculate $EngagementValue_{i}=(1-w(d_{i}))\cdot y^{E}_{i}$
     if $MonetizationValue_{i}\leq EngagementValue_{i}$ then
        if $\min(|\Delta y^{E1}_{i}|,\dots,|\Delta y^{Ek}_{i}|)\geq\max(|\Delta y^{E1}_{i}|,\dots,|\Delta y^{Ek}_{i}|)$ then
           Recommend “Boost Engagement”
        else
           Recommend “Promote Upsell”
        end if
     else
        if $y^{U}_{i}>0$ then
           Recommend “Promote Upsell”
        else
           Recommend “Prevent Churn”
        end if
     end if
  end for
4.1.5. Output
The optimization layer outputs the following for each account (Table 1 & 2):
Table 1. Optimization Layer Outputs - Account Details
Account ID Rep ID gRank111gRank :account ranking across all reps based on objective values rRank222rRank :account ranking within the assigned rep based on objective values
101 R1 1 1
102 R1 2 2
103 R2 3 1
Table 2. Optimization Layer Outputs - Actions and Metrics
Account ID Action $U_{i}$ $E_{i}$
101 Promote Upsell 5000 15
102 Boost Engagement 2000 30
103 Prevent Churn -1000 5
4.2. Contextual Bandit
After constrained optimization, the recommendations are processed by a contextual bandit layer, which directly interacts with the final serving layer. This layer is integral to the system, leveraging recommendations from upstream modules, such as Monetization Upsell, Monetization Churn, and Low Engagement mapped at the sales representative level, to optimize these recommendations further using user actions and finally sending them to the serving layer. By combining contextual features of recommendations with user actions/feedback, the contextual bandit iteratively refines its policy, dynamically balancing exploration (to improve understanding of less tried recommendations) and exploitation (to serve the most effective recommendations).
Each sales representative is assigned a set of accounts determined by the constraint optimization step. The contextual bandit selects an action from the following predefined set:
  • $a^{B}$ (Boost Engagement)
  • $a^{C}$ (Prevent Churn)
  • $a^{U}$ (Promote Upsell)
These actions are encoded as contextual features within the model to refine decision-making. Over time, the bandit learns which action maximizes cumulative engagement and business value.
4.2.1. Reward Mapping and Objective Function
The contextual bandit framework leverages user feedback to iteratively refine action selection. The system assigns a reward to each action based on observed user responses:
  • $\mathcal{A}=\{a^{B},a^{C},a^{U}\}$ is the set of actions.
  • Reward from user feedback is defined as,
    $y=\begin{cases}+1&\text{if user feedback = \text{DEEP\_LINK\_CLICKED},}\\ -1&\text{if user feedback = \text{NOTIFICATION\_DISMISSED},}\\ 0&\text{if user feedback = \text{NO\_CLICK}.}\end{cases}$
The overarching goal of the contextual bandit layer is to identify the optimal policy $\pi^{*}$ that maximizes cumulative rewards over time. Formally, this objective can be expressed as:
$\pi^{*}=\arg\max_{\pi}\mathbb{E}\left[\sum_{t=1}^{T}y^{\pi}_{t}(a_{t},x_{t})\right],$
where:
  • $a_{t}\in\mathcal{A}$ is the action taken at time $t$.
  • $x_{t}$ represents contextual features at time $t$.
  • $y^{\pi}_{t}(a_{t},x_{t})$ denotes the reward from user feedback by following an action policy $\pi$.
This formulation ensures that the system progressively converges to a policy that aligns recommendations with user preferences and optimizes engagement over time.
4.2.2. Contextual Feature Representation
The contextual bandit layer incorporates a rich set of features to enhance decision-making. The contextual feature vector is defined as:
$x_{t}=[x^{A}_{t},x^{S}_{t},x^{R}_{t}],$
where:
  • $x^{A}_{t}$ (Account Features): Includes attributes such as account size, industry, engagement level.
  • $x^{S}_{t}$ (Sales Representative Features): Encapsulates experience, historical success rate, and past interactions.
  • $x^{R}_{t}$ (Recommendation Features): Captures details like type of recommendation and predicted impact.
These features are extracted using:
$x_{\text{numerical}}=\begin{bmatrix}\text{TimeSinceLastAlert}\\ \text{PreviousAlertCount}\end{bmatrix},$
with one-hot encoded categorical representations:
[…]
The resulting vector $x_{t}$ ensures that the model evaluates each action within a context-sensitive framework.
4.2.3. Policy Optimization
Policy optimization is achieved by employing Neural Bandit (Dai et al., 2022; Zhou et al., 2020). Each action’s reward distribution is modeled as:
$y_{a}=h(x_{t},a_{t})+\epsilon$
where:
  • $h(x_{t},a_{t})$ is a unkown functional mapping from feature to the expected reward.
  • $\epsilon$ is a sub-Gaussian noise.
We approximate the unknown function $h(\cdot)$ with a neural network $f_{\theta}(\cdot)$. The algorithm iterate over these steps:
  1. (1)
    Compute contextual feature representation $x_{t}$.
  2. (2)
    Predict reward for each action $a\in\mathcal{A}$ using the neural network:
    $\hat{y}_{a}=f_{\theta}(x_{t},a)+\delta.$
  3. (3)
    Select the action that maximizes the predicted reward:
    $a_{t}=\arg\max_{a\in\mathcal{A}}\hat{y}_{a}.$
  4. (4)
    Observe user feedback and update network parameters $\theta$ via backpropagation.
Here, value of $\delta$ depends on a key variance quantity $\sigma$,
$\sigma=\sqrt{\nabla f(\cdot)^{\top}H^{-1}\nabla f(\cdot)}$
where:
  • $\nabla f(\cdot)$ is the gradient of the neural network output with respect to input features.
  • $H^{-1}$ is the inverse Hessian matrix used to capture model uncertainty.
In Thompson Sampling, we sample $\delta$ from Gaussian distribution, $\delta\sim\mathcal{N}(0,\beta^{2}\sigma^{2})$, and in Upper Confidence Bound, we calculate $\delta=\gamma\sigma$, where both $\beta$ and $\gamma$ are tuning parameter controlling the balancing of exploration and exploitation.
Through these steps, the contextual bandit ensures that engine evolve to serve the most relevant recommendation to the user effectively while optimizing long-term engagement.
5. Serving Layer
The Serving Layer in the System serves the recommendation to sales functions and further transforms predictive scores and features into actionable insights for sales representatives. By employing a structured and interpretable methodology, it ensures recommendations are clear, precise, and aligned with business objectives. Our framework has been designed to support template based explanations and template-free explanations powered by GAI.
5.1. Template based explanations
5.1.1. Formulating the Explanation Layer
The Explanation Layer is built on three primary components: feature mapping, threshold definitions, and templates. It translates model outputs into human-readable insights through:
  1. (1)
    Mapping feature names to human-understandable expressions.
  2. (2)
    Defining thresholds for key features and model combinations.
  3. (3)
    Generating recommendations based on these mappings and thresholds.
5.1.2. Feature Mapping
A manual mapping translates feature names into comprehensible language. (Example: Table 7 in the Appendix):
5.1.3. Threshold Definitions
Thresholds are defined for each feature and model combination to determine when a specific recommendation is triggered. (Example: Table 8 in the Appendix)
5.1.4. Templates for Recommendations
Templates are designed to produce tailored recommendations for each account. These recommendations are categorized into key alert types and include structured explanations for feature values that pass thresholds. Examples include:
  • Low Engagement: RTCD = $d$. We recommend reaching out to the client to understand the low engagement in ¡product¿. The current usage is $y$ and we are predicting a drop to $\Delta y$ over the next month.
  • Upsell Flag: RTCD = $d$. We recommend exploring add-on opportunities with this customer as we predict a near-term upsell opportunity worth $\Delta y$.
  • Churn Flag: RTCD = $d$. We recommend connecting with customers to assess churn risks.
5.2. Template-free explanations with GAI
To enhance the explanation layers and streamline the on-boarding process, we leverage LLM to generate feature explanations for each feature name, and integrate Generative AI (GAI) with sales productivity tools to automate narrative generation and feature categorization. Multiple instance-level explanation algorithms (e.g., CLIME, TE2Rules, Integrated Gradients) are incorporated for calculating the importance of instance-level features (Figure  2). Using an iterative approach, we specify the necessary steps to complete a task and provide a sample output format for GAI to learn. By integrating feature names and explanations along with instance-level feature importance scores into the LLM, we generate detailed instance-level explanations  [^2] For detailed steps, please refer to Appendix A.5.. This includes insights into what changes in each feature might indicate, offering a comprehensive understanding of the data’s impact on model recommendations (Figure  3).
Figure 2. GAI Based Instance Level Explanation Generation Process
Figure 3. GAI Based Instance Level Explanation Generation Process
The framework can also integrate multiple components to generate insights and provide conversational experience for user in agentic AI framework (Figure  4). The Content Component processes input data, including tabular, text, and image data, through an Input Data Module. This data is then used in the Insights Module to generate sales funnel-specific, sales call, product and engagement insights, and other relevant insights, which are standardized and normalized. The Agent Component interacts with users, utilizing these insights and generating conversation logs. These logs, including labeled prompt-response pairs, are used in the Fine-Tune module to improve the agentic AI system.
Figure 4. GAI Based Instance Level Explanation Generation Process
6. Experiments and Results
We conduct experiments to evaluate our framework using both offline and causal inference. Our evaluation has two main objectives:
  1. (1)
    Evaluate the individual performance of the layers and components using offline data
  2. (2)
    Analyze the business impact of our solution by deploying to a live system and collecting user feedback
6.1. Offline Analysis and Validation
6.1.1. Offline Evaluation Dataset
Training data for our offline evaluation consists of active SME customers, with closed opportunities, during the data collection period. These closed opportunities were either won or disengaged. Our features included customer engagement metrics, historical bookings, product usage, company firmographics, and LinkedIn-only data like hiring trends.
6.1.2. Prediction Layer Implementation
The predictive models utilized in the prediction layer were trained and evaluated on historical data consisting of product engagement, previous customer transactions, and Go-To-Market (GTM) interaction metrics. GTM interaction metrics include sales outreach, sales offers, and other incentives used to retain or grow customer accounts over time. Other features include company firmographics and LinkedIn’s proprietary customer data like hiring trends, among others. The training data are refreshed regularly using LinkedIn proprietary orchestration platform which ensures the prediction layer models are updated and capture concept drift, covariate shift, and label drift over time. The prediction layer also handles data quality concerns, such as noisy data and sparse historical records. Finally, the prediction layer utilizes a robust pipeline for training the model used by the monetization uplift and customer engagement forecast models. Our robust pipeline prevents overfitting the model using cross-validation and utilizes data monitoring tools to detect drift in model performance over time.
6.1.3. Uplift Model Evaluation
The uplift model estimates the Individual Treatment Effect of sales actions on monetization metrics for each customer. We compared four Conditional Average Treatment Effects (CATE) estimators for the uplift model: S-learner, T-learner, X-learner and DR Learner (Künzel et al., 2019; Robins et al., 1994; EconML, 2019) (Details in Appendix  A.1).
Evaluating uplift models can be challenging due to the lack of ground truth for counterfactual outcomes, which makes it difficult to find a loss measure for each observation. We use the uplift deciles (Gutierrez and Gérardy, 2017) for aggregated measures such as uplift bins or uplift curves(Figure  5). Our evaluation also focused on internal business considerations such as prediction stability and score interpretability as secondary metrics.
Figure 5. Uplift Deciles for Baseline Models
While the results from the S-learner and X-learner models mimic the expected ITE distribution, we selected the T-learner for our uplift modeling due to the consistency and stability of the decile bins and ITE distributions over multiple iterations.
6.1.4. Customer Engagement Models Evaluation
Customer engagement metrics are continuous values, that measure monthly product adoption and product utilization for each customer. We denote these metrics as $pa$ and $pu$ respectively. We evaluated model accuracy using Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE). Table 3 summarizes the prediction accuracy results for our models $f_{pu}$ and $f_{pa}$.
Table 3. Forecasting models prediction accuracy metrics
Model Score Range MAE RMSE
$f_{pu}$ 0- 100 7.581 10.775
$f_{pa}$ 0 - 1 0.203 0.291
6.1.5. Ablation analysis for Constrained Optimization
We conducted an ablation analysis of the optimization layer by systematically removing or modifying key components—such as the weighting function, capacity constraints, and recommendation rules. Using six months of backtesting data, we compared the performance of the Full Model (with all components intact) to simplified models across key metrics, including upsell precision, churn precision, churn recovery precision, low engagement precision, total bookings rank, and constraint feasibility (Details in Appendix  A.2).
The Full Model consistently outperformed the ablated models, demonstrating the importance of each component:
6.1.6. Validation for Contextual Bandit
To evaluate the contextual bandit layer’s performance, we conducted a comprehensive offline analysis. This analysis aimed to validate the bandit’s ability to optimize recommendations based on contextual information and observed rewards. The key areas of analysis include:
  • Feedback Distribution: This analysis examines how frequently each feedback, i.e., action from sales representatives, (DEEP_LINK_CLICKED, NOTIFICATION_DISMISSED, NO_CLICK) was selected during the simulations. The distribution highlights the bandit’s prioritization of high-reward feedback (Figure  7 in Appendix).
  • Feedback Trends Over Time: Temporal trends in Feedback percentages provide insights into the bandit’s progression from exploration to exploitation, i.e., focusing on actions that leads high-reward feedback (Figure  6).
    Figure 6. Feedback Trend Over Time for Thompson Sampling
For the bandit optimization, both Neural Thompson Sampling and Neural UCB approaches were evaluated to compare their effectiveness in terms of policy convergence, cumulative rewards, and their ability to balance exploration and exploitation. The cumulative rewards for both strategies are comparable over multiple iterations (Figure 8 in Appendix). The current system adopts Thompson Sampling due to its low computational overhead per iteration leading to low latency while serving.
6.2. Roll-out and Launch Details
CPOG’s modular and flexible design enables tailored solutions for various customer segments and regional needs. We launched a test within LinkedIn, targeting 25% of sales reps focused on SME customers in North America. Sales reps received personalized recommendations which are delivered in near real-time through CRM applications, seamlessly integrating with sales workflows. Feedback on recommendation relevance is collected to continuously improve future suggestions. Our solution received high user satisfaction among sales reps (details in Appendix  A.3) and delivered significant business impact (+59.96% lift on primary metrics) with minimal disruption to sales workflows and quota achievement.
6.2.1. Challenges to A/B Test
While A/B test is the best practice for impact estimation, several challenges must be considered in our scenario
  1. (1)
    Opportunity Costs: A partial rollout could limit solution availability, resulting in missed opportunities despite its potential to alleviate sales representatives’ pain points with minimal disruption.
  2. (2)
    Business Urgency: Achieving revenue goals amidst macroeconomic pressures necessitates leveraging the solution’s timely insights promptly.
  3. (3)
    Selection Bias: Randomization in treatment group selection may be infeasible, introducing potential selection bias.
  4. (4)
    Quota Equity: Disparities in quotas arising from the test must be mitigated through appropriate quota adjustments during interim sales planning cycles.
  5. (5)
    Statistical Power: Testing with a small group may lack sufficient power to produce statistically significant results.
Given the challenges of conducting a randomized A/B test and the subsequent decision to implement a full rollout after the initial ramp, we implemented an observational causal study to estimate the business impact.
6.2.2. Causal Inference
The treatment and control groups were defined as follows:
  • Treatment Group: Sales representatives in North America who had access to the recommendations
  • Control Group: Sales representatives outside North America without access to the recommendations
The units in our treatment group are ¡rep, account¿ pairs because recommendations are generated for only a subset of accounts within each sales reps’ books. We focus exclusively on relevant booking opportunities [^3] Opportunities originating from accounts with valid recommendations, generated within a specified time period, and followed up by sales outreach which makes our treatment period time-dependent. To account for irregular and sparse timing of treatments across multiple periods, we utilized Difference-in-Difference (DiD) (Ashenfelter and Card, 1984) methodology for our causal analysis. DiD estimates the Average Treatment Effect on the Treated (ATT) - mean difference in actual outcomes for the treated units and their counterfactual outcomes.
$ATT=\mathbb{E}[Y(a)-Y(0)|a=1]$
We report the Relative Treatment Effect (RTE) of our ATT estimate - which gives a better intuition of the treatment effect relative to the baseline or control group.
$RTE=ATT/D$
Where $D=\mathbb{E}[Y(a)|a=1]-\mathbb{E}[Y(a)-Y(0)|a=1]$
6.2.3. Assumption Validation
To validate the parallel trends assumption, we conducted a time-based A/A test (also known as pre-testing). In this approach, we used the same treatment and control groups as in our main analysis but ran placebo Difference-in-Differences tests at various points in the pre-treatment period. For each test, we treated an earlier time point as if it were the treatment period and evaluated whether the estimated treatment effect was close to zero. Since no intervention had actually occurred during these pre-periods, finding no significant differences supports the assumption that the treatment and control groups were following similar trends before the intervention. We also conditioned on important covariates like account size, engagement level, and past sales activity to ensure balance.
6.2.4. Experiment Setup
Our online experiments consists of two time periods: (1) pre-intervention - during the second half of 2023 and (2) post-intervention - during the first half of 2024. Our control group consists of candidate accounts outside North America and were matched to the treatment group using Coarsened Exact Matching (CEM) (Iacus et al., 2012). We used a proxy metric, Net Ratio, as the primary response variable for our test period and define Net Ratio as follows:
[…]
6.2.5. Results
The causal inference analysis demonstrated a significant improvement in the primary metric, Net Ratio (NR), when compared to the legacy system. As defined in Equation 7, Net Ratio captures account-level monetization effectiveness.
Table 4. Difference-in-Difference Results on Net Ratio (NR)
Method Relative Treatment Effect on NR p-value
Legacy N/A N/A
CPOG 59.96% 0.0178
7. Conclusion
In this paper, we introduce the Causal Predictive Optimization and Generation (CPOG) framework, a novel approach to sales optimization and business AI. Through offline analysis, online testing and deployment, we demonstrate CPOG’s effectiveness in optimizing sales strategies across multiple metrics. In particular, we highlight the importance of using uplift models to prioritize sales actions, the benefits of constrained optimization in balancing multiple objectives, and the value of providing sales representatives with clear, explainable and actionable recommendations. We give guidelines on how to apply CPOG framework and clear examples of how we implemented it at LinkedIn, along with results and potential measurement approaches in complex applications where A/B testing is not feasible. The framework and the learnings in deployments are broadly applicable to the field, particular B2B and SaaS businesses.
Appendix A Additional Offline Analysis
A.1. Comparison for Uplift Modeling
  • S-learner (Künzel et al., 2019) uses a single estimator that includes all features and treatment indicators without giving the treatment indicator a special role. While this simplified structure makes the S-learner easy to implement, not assigning the indicator a special role means the model may or may not use the indicator feature for modeling.
  • T-learner (Künzel et al., 2019) or Two-learner is the most common meta-algorithm for estimating heterogeneous treatment effects by building two regression models trained on the control and treatment groups separately. The treatment effect is estimated as the difference between the predictions from the regression models. T-learners can be biased if either regression model is trained on insufficient data
  • X-learner (Künzel et al., 2019) extends the concept of T-learners by using each observation in the training set in an ”X”-like shape. This allows for estimating the CATE by regressing the difference of the ITEs on the covariates. X-learner improves on the limitations of other meta-learners - it is more effective for instances where the training data volume for one outcome is greater than the other. The need for large training data is also a limitation of X-learners.
  • DR Learner (Robins et al., 1994) The DR learner estimates treatment effects under the assumption that all confounders are observed by breaking the problem into two predictive tasks: predicting the outcome from treatment and controls, and predicting treatment from controls. It then combines these models in a final stage to estimate heterogeneous treatment effects.
A.2. Details of Ablation Analysis for Optimization
A.2.1. Objective
To evaluate the effectiveness of each primary component in the optimization framework, including feature mappings, constraints, and recommendation rules. The objective is to quantify their individual contributions to monetization outcomes and overall model effectiveness.
A.2.2. Experimental Setup
We performed simulations under various configurations by systematically removing or modifying key components of the optimization layer. The performance was evaluated using six months of backtesting data to measure the impact of each component.
  • Full Model (Baseline): Includes all components—weighted objective function, capacity constraints, RTCD prioritization, and eligibility filters.
  • Ablated Models:
    1. (1)
      Model A (No Weighting Function): Removed the weighting function $w(d_{i})$, treating monetization uplift and engagement improvement equally across all accounts.
    2. (2)
      Model B (Relaxed Capacity Constraints): Removed capacity constraints, allowing sales representatives to be assigned any number of accounts.
    3. (3)
      Model C (Simplified Recommendation Rules): Simplified the recommendation rules to always prioritize monetization uplift.
A.2.3. Metrics
  1. (1)
    Upsell Precision ($P_{\text{ups}}$): The percentage of accounts recommended for upsell that successfully closed as upsell opportunities.
  2. (2)
    Churn Precision ($P_{\text{ch}}$): The percentage of accounts recommended as churn that actually churned without sales outreach.
  3. (3)
    Churn Recovery Precision ($P_{\text{rec}}$): The percentage of accounts recommended as churn that were successfully retained through sales outreach.
  4. (4)
    Low Engagement Precision ($P_{\text{low}}$): The percentage of accounts recommended as low engagement that churned without sales outreach.
  5. (5)
    Total Bookings Rank ($B_{\text{total}}$): The total revenue target achieved (RTA) across all account categories ranked in descending order.
  6. (6)
    Model Feasibility (Constraints Met): The percentage of capacity and assignment constraints satisfied during optimization.
A.2.4. Results
Table 7 and 8 are results from backtesting data.
Table 5. Ablation Analysis Results - Model Performance
Model $P_{\text{ups}}$ $P_{\text{ch}}$ $P_{\text{rec}}$ $P_{\text{low}}$
Full 60% 55% 50% 40%
Model A 56% 45% 42% 30%
Model B 35% 26% 37% 25%
Model C 52% 43% 38% 48%
Table 6. Ablation Analysis Results - Ranking and Constraints
Model $B_{\text{total}}$ Rank % Constraints Met
Full Model 2 100%
Model A (No Weighting) 3 100%
Model B (No Constraints) 1 216%
Model C (Simplified Rules) 4 100%
The ablation analysis highlights the critical contributions of each component in the optimization layer to its overall effectiveness. The Full Model’s superior performance demonstrates the necessity of incorporating all key features:
  • Weighting Function: Essential for dynamically adjusting priorities between monetization uplift and engagement improvement based on account proximity to RTCD.
  • Capacity Constraints: Crucial for ensuring practical and equitable assignment of accounts while maintaining feasibility across operational workflows.
  • Nuanced Recommendation Rules: Integral to balancing diverse objectives, such as upsell, churn recovery, and engagement improvement, to achieve holistic optimization.
While simplified models (e.g., Model B and Model C) may yield partial gains in specific metrics, they compromise overall balance and feasibility. Therefore, the Full Model provides the most robust solution for optimizing sales strategies, ensuring both short-term and long-term business objectives are met effectively.
A.3. Survey on Explanation Layer
We conducted a survey on the explainability of our recommendations across 142+ accounts, achieving an overall 86% satisfaction rate, with strong positives on comprehensiveness and clarity.
Key feedback highlights:
  • Clear and actionable insights build trust in recommendations and uncover new opportunities.
  • Consolidated insights save time, reducing the need to gather data from multiple sources.
  • Explanations provide directional guidance, helping reps take right next steps in customer engagement.
  • Sales reps’ top signals for consideration align with the top important features from the explanation layer, reinforcing trust in the recommendation logic.
A.4. Explanation tables
Table 7. Feature Name Mapping
Feature Name Expression
producta_l1m Avg. product A usage for last month
productb_l2m Avg. product B usage across last 2 months
Table 8. Basic Information and Thresholds
Feature Name Model Threshold
producta_l1m Treatment Model ¿0
producta_l1m Control Model ¡0
Figure 7. Feedback Distribution for Thompson Sampling
Figure 8. Comparison of Cumulative Rewards Between Thompson Sampling and UCB.
A.5. GAI-Based Instance-Level Explanation Generation Process
This appendix outlines the multi-stage process for generating interpretable, instance-level explanations by combining algorithmic feature importance with semantic grouping.
A.5.1. Step 1: Feature Grouping using LLM
Given a list of raw feature names, we use LLM to group them semantically based on suffix patterns and domain expertise:
  • super_name: Meaning of features.
  • ultra_name: Higher-level category that unifies related features across timeframes.
Mapping Rules: Internal shorthand and acronyms are expanded (e.g., metric prefixes) and grouped accordingly. Related engagement metrics (e.g., Metric A, Metric B) are clustered under broader semantic categories.
Example Output:
Feature_name super_name ultra_name
MetricA_l1m Metric A in the last month Metric A
MetricA_l2m_l1m Metric A in the last month Metric A
MetricA_l3m Metric A in the last 3 months Metric A
MetricB_l1m Metric B in the last month Metric B
Table 9. Abstracted Feature Grouping Example
A.5.2. Step 2: Combining Feature Importance with Metadata
We extract instance-level feature importance from the instance-level feature importance algorithm and join it with grouped feature metadata. This includes:
  1. (1)
    Extracting feature importance scores from instance-level feature importance algorithm.
  2. (2)
    Joining with the semantic feature grouping table (containing super_name and ultra_name).
  3. (3)
    Enriching with feature values per account.
Output Schema:
Weight Feature_name Value
0.072 MetricC_l12m 24.0
Table 10. Step 2 output table 1
Super_name Ultra_name customer_urn
Metric C in the last 12 months Metric C urn:li:customer:…
Table 11. Step 2 output table 2
A.5.3. Step 3: Explanation Generation using LLM
We use LLM to generate ranked natural language explanations based on weighted features. The generation logic:
  • One insight per distinct Ultra_name, ordered by descending feature weight.
  • Combine narratives for related metrics (e.g., all hiring-related features grouped into a single insight).
Example Output:
This account is likely a good candidate to Promote Upsell. Its likelihood is driven by:
1. In the past 12 months, Metric C increased from 18 to 24 (+33%).This shows a high level of engagement and receptivity from the target audience.
2. Metric A in the last month increased from 78% to 85% (+7%), which reflects a high level of brand awareness and interest among potential applicants.